OpenStudy (anonymous):

y varies jointly with p and q. When p = 5 and q = 2, y = 80. Find y when p = 4 and q = 5. y = 20 y = 120 y = 160 y = 240

OpenStudy (anonymous):

\[y=pq(k)\]

OpenStudy (anonymous):

80 = 5(2)(k)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

now find y

OpenStudy (anonymous):

80 = 10k

OpenStudy (anonymous):

find y when p=4 q=5 and the k you got k=8

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

I'm stuck :/

OpenStudy (anonymous):

80=10(k) ---- you will find k here

OpenStudy (anonymous):

which is 8 right?

OpenStudy (anonymous):

right...

OpenStudy (anonymous):

now apply the equation again y=pq(k)

OpenStudy (anonymous):

but now using p = 4, q = 5 and k =8

OpenStudy (anonymous):

you should get the Ynew = (4)(5)(8)

OpenStudy (anonymous):

y / (pq) = k 80 / (5•2) = k 8 = k y / (4•5) = 8 y / 20 = 8 y = 8 • 20 y = 160?

OpenStudy (anonymous):

nice. you got nit right

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no problem, goodluck solving math problems.